Blue Spruce Inc. bought a business that is expected to give a 26% annual rate of return on the investment. Of the total amount paid for the business, $76,300 was deemed to be goodwill, and the rest was attributed to the identifiable net assets. Blue Spruce Inc. estimated that the annual future earnings of the new business would be equal to the average ordinary earnings per year of the business over the past three years. The total net income over the past three years was $374,000. This amount included a loss on discontinued operations of $22,800 in one year and an unusual and non-recurring gain of $93,600 in one of the other two years.

Calculate the fair value of the identifiable net assets that Blue Spruce Inc. purchased in this transaction.
Answer #1

Total income before adjustments for 3 years is $37400. After adjustment of abnormal loss & non-recurring gain, Net Total income for 3 years is 374000+22800-93600 = $303200.

Average of 3 year's income = $303200/3 = $101067, which is 26% return on total investments.

Hence Total Investments = $10106/26% = $388718.

Now, Total Investments = Goodwill + identifiable net assets,

Therefore, identifiable net assets = Total Investments - Goodwill = $388718 - $76300 = $312418.